Solve the equation by the​ zero-factor property.
s^2-12s+35=0

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

s^2-12s+35=0
(s-7)(s-5)=0
s=7,5

Answer by ikleyn(52776)

.
s^2 - 12s + 35 = 0.

Factor left side:

(s-5)*(s-7) = 0.

The roots are s= 5 and/or s= 7.